OpenStudy (trisarahtops):

The end behavior of f(x) = 9+2x^3 / x^4-81 most closely matches which of the following? y = 0 y = 2 y = 9 y = −9

OpenStudy (solomonzelman):

Evaluate f(10) f(100) f(1000) (use a calculator) and tell me when happens to the function as x grows.

OpenStudy (trisarahtops):

okay gimme a sec

OpenStudy (trisarahtops):

f(10)= .2025

OpenStudy (trisarahtops):

f(100)= .02000

OpenStudy (trisarahtops):

it gets smaller @SolomonZelman

OpenStudy (solomonzelman):

Yes, and it would continue getting smaller and smaller....

OpenStudy (solomonzelman):

it actually goes to zero...

OpenStudy (solomonzelman):

Do you want to leave with that, or do you want to verify that this is so?

OpenStudy (trisarahtops):

alright so zero is the answer. I've got it thanks ;)

OpenStudy (solomonzelman):

For very large values of x (like x=1000 and bigger) \(\large\color{#000000 }{ \displaystyle 2x^3>9>0 }\) and (also for very large x-values); \(\large\color{#000000 }{ \displaystyle \frac{1}{2}x^4<x^4-81<x^4 }\) And the following would be also true: \(\large\color{#000000 }{ \displaystyle \frac{2x^3}{x^4}<\frac{9+2x^3}{x^4-81} <\frac{2x^3+2x^3}{x^4-\frac{1}{2}x^4} }\)

OpenStudy (solomonzelman):

Do you agree with the last statement? (left side => smaller numerator, bigger denominator right side => bigger numerator, smaller denominator so, left < middle < right)

OpenStudy (solomonzelman):

What will this do? We can simplify both right and left sides and show that they surely go to zero, and therefore the middle (our function would also go to 0.

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \frac{2x^3}{x^4}<\frac{9+2x^3}{x^4-81} <\frac{2x^3+2x^3}{x^4-\frac{1}{2}x^4} }\) \(\large\color{#000000 }{ \displaystyle \frac{2}{x}<\frac{9+2x^3}{x^4-81} <\frac{24x^3}{\frac{1}{2}x^4} }\) \(\large\color{#000000 }{ \displaystyle \frac{2}{x}<\frac{9+2x^3}{x^4-81} <\frac{48}{x} }\)

OpenStudy (solomonzelman):

and left and right surely go to zero.